OpenStudy (anonymous):

radius =16 angle of the sector =60 what is the arc length of the circle

OpenStudy (anonymous):

60 Degrees right?

OpenStudy (anonymous):

60 degree is the angle of the sector i need the arc length of the sector

OpenStudy (anonymous):

Ok. The Formula you need is: Angle * (Pi/180) * Radius This Formula is when you use the Angle in Degrees. When you use Radians just forget the (Pi/180) part.

OpenStudy (anonymous):

16 *\[16\times (\pi \div180)\times60\]

OpenStudy (anonymous):

Wait a little minute: Although it doesn't really matter, The 60 and 16 should have been written in each others place as I wrote Angle * (Pi/180) * Radius

OpenStudy (anonymous):

The answer is: 16.755 (to 3 decimal places) or\[16/3\pi\]